Compute NOPAT Selected income statement information for 2018 is presented below for Home Depot Inc. and Lowes Companies in P i need help with 1,2,3
Answer #1

Solution:

a) Here we have 2 companies , Home Depot and Lowe.

1. Net Operating Profit (NOPAT) IS

= Net Operating Profit Before Tax - Tax expense

Net Operating Profit of 2 companies are

Home Depot :  

= 15841 - 3504 = $12337

Lowe's  :

= 4098 - 1080 = $3018

2. Return on Net Operating Assets (RNOA)

= Net Operating Profit / Average Net Operating Assets

Home Depot

= 12337/ 25721 = 47.9%

Lowe's

= 3018/ 20326 = 14.8%

3. Net Operating Profit Margin (NOPM) is

= Net Operating Profit / Sales

Home Depot

= 12337/ 108203 = 11.4%

Lowe's

= 3018/ 72735 = 4.14%

4. Net Operating Asset Turnover (NOAT) is

= Sales/ Average Net Operating Asset

Home Depot

=108203/ 25721 = 4.20

Lowe's

= 72735/ 20326 = 3.57

b)

1. More profitable ( in $ ) is Home Depot. Because Home Depot Net Operating Profit Margin is higher than Lowe's net operating profit margin.. Home depot 11.4% and Lowe's 4.14%.

2. Home depot produces the higher profit margin (%) .. Home depot 11.4% , Lowe's 4.14%

   3. Home depot uses its Net Operating Asset ( NOA) more efficiently. Because Home depot Return on Asset is higher than Lowe's. Home depot 47.9% , Lowe's 14.8%

4. Home depot produces the higher return on Net Operating Asset.It indicates that management is maximizing the use of the company's assets.